OpenStudy (anonymous):

You play a game in which two dice are rolled. If a sum of 7 appears, you win $10; otherwise, you lose $2.00. If you intend to play this game for a long time, should you expect to make money, lose money, or come out about even? Explain.

OpenStudy (anonymous):

find the expected value

OpenStudy (anonymous):

probability a 7 appears is \[\frac{1}{6}\] probability it doesn't appear is \[\frac{5}{6}\] expected value is \[10\times \frac{1}{6}-2\times \frac{5}{6}=\frac{10-10}{6}=0\] so gaiem is "fair"

OpenStudy (anonymous):

*game expected value is zero, you don't expect to win or lose in the long run

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@bmar1209 I hope you know where the 1/6 comes from!

OpenStudy (anonymous):

Great! I'm happy to walk away now :)